OpenStudy (anonymous):

what are the minimum and maximum values of x+sin^2 x with the domain of pi/6 to 5pi/6?

OpenStudy (perl):

f(x) = x + sin^2(x) to find max/min, find the critical points, then use second derivative test

OpenStudy (perl):

is this a closed domain?

OpenStudy (anonymous):

its greater than or equal to

OpenStudy (perl):

pi/6 <= x <= 5pi/6 ?

OpenStudy (anonymous):

yes

OpenStudy (perl):

ok then we can find the absolute max and min (which means we need to test the endpoints as well)

OpenStudy (perl):

go ahead and find the derivative, and set it equal to zero

OpenStudy (anonymous):

i found that the first derivative 1+2sincos

OpenStudy (perl):

ok, and 2sinx cosx is equal to sin(2x) by a trig identity

OpenStudy (perl):

f ' (x) = 0 1 + 2 sin x cos x = 0 1 + sin(2x) = 0 sin(2x) = -1

OpenStudy (anonymous):

what is sin(2x)=-1?

OpenStudy (loser66):

that is for f' =0 when sin (2x) =-1, solve for x is the next step

OpenStudy (perl):

we can take inverse sine

OpenStudy (anonymous):

which is cos

OpenStudy (perl):

not quite. i meant like arcsine

OpenStudy (perl):

sin(2x)=-1 <=> 2x = sin^-1 ( 2x )

OpenStudy (loser66):

 so, when sin (something ) =-1? what is that "something"?

OpenStudy (loser66):

3pi/2, right? that is sin (3pi/2) =-1, right?

OpenStudy (anonymous):

right

OpenStudy (loser66):

now, our "something" is 2x , so that 2x = 3pi/2, right? then x = ???

OpenStudy (anonymous):

3pi?

OpenStudy (loser66):

hey, 2x = 3pi/2, x = 3pi/2/2= 3pi/4 hahaha.... take a rest, friend.

OpenStudy (loser66):

no??

OpenStudy (loser66):